OpenStudy (anonymous):

factor completely (2y-3)^2-8(2y-3)+16 answers : a)(2y-4)^2 b)(2y-3)^2+16 c)[(2y-3)+4]^2 d)(2y-7)^2 e) (2y-1)^2

OpenStudy (anonymous):

let u= 2y-3

OpenStudy (anonymous):

u^2-8u+16 =0 (u-4)^2=0

OpenStudy (anonymous):

^ shouldnt have put equals zero up there, but same thing

OpenStudy (anonymous):

( 2y-3-4)^2 =(2y-7)^2

OpenStudy (anonymous):

so d?